OpenStudy (fibonaccichick666):

Define f(x) ={\(x^2+2x^2sin\frac{1}{x}~~~x\not=0\);\(0~~x=0\)} Show that f'(0)>0.

OpenStudy (fibonaccichick666):

@ganeshie8 , I am confused because the derivative will be 0 at 0, right?

OpenStudy (fibonaccichick666):

by squeeze it is zero

OpenStudy (solomonzelman):

derive and plug in zero for x into the derivative(?)

OpenStudy (fibonaccichick666):

yea, I get 0 is the issue

OpenStudy (solomonzelman):

zero for f'(x) (before plugging in zero for x into the derivative) ?!

OpenStudy (solomonzelman):

I am sure you can derive x^2, and the 2x^2sin(1/x) by the product rule, with a chain for sin(1/x).

OpenStudy (fibonaccichick666):

omg... ok ignore me now.

OpenStudy (fibonaccichick666):

give me a minute, i was still doing the last problem's limit

OpenStudy (solomonzelman):

sure

OpenStudy (solomonzelman):

maybe you need help on the last problem (whatever it is) ?

OpenStudy (fibonaccichick666):

so for the deriv. I have 2x+4xsin(1/x)+-2cos(1/x) which gives me a dilemma at x=0

OpenStudy (solomonzelman):

your derivative will not be defined at x=0.

OpenStudy (solomonzelman):

but you are not getting the correct derivative.

OpenStudy (fibonaccichick666):

?

ganeshie8 (ganeshie8):

\[f'(0) = \lim\limits_{x\to 0} \dfrac{f(x) - f(0)}{x-0} = \lim\limits_{x\to 0}~ x + 2x\sin(1/x) \]

OpenStudy (solomonzelman):

f(x) = x^2+2x^2 sin(1/x) f '(x) = 2x + 4x sin(1/x) + 2x^2 cos(1/x) * d/dx(1/x)

OpenStudy (solomonzelman):

f '(x) = 2x + 4x sin(1/x) + 2x^2 cos(1/x) * (-1/x^2)

ganeshie8 (ganeshie8):

you want to apply squeeze thm for evaluating that limit ? let me grab a paper in between

OpenStudy (solomonzelman):

Chain for the inner function of the sine.

OpenStudy (fibonaccichick666):

I fail to see the difference between our answers

OpenStudy (fibonaccichick666):

but anyways, thanks ganesh, I think I'm gonna go from the def. Good idea

OpenStudy (solomonzelman):

I will not disturb, but I'll only say that when you tried to derive the second function, the derivative of sin(1/x) is not just cos(1/x), it is cos(1/x) times the derivative of 1/x. (CHAIN RULE)

OpenStudy (fibonaccichick666):

@ganeshie8 , I'm getting 0 after squeeze though

ganeshie8 (ganeshie8):

so it doesn't really help in showing what u want to show.. hmm

OpenStudy (solomonzelman):

Do your definition, as ganeshie says.

OpenStudy (fibonaccichick666):

yes, and \(-1/x^2 *2x^2=-2\)

OpenStudy (fibonaccichick666):

I simplified solomon

OpenStudy (solomonzelman):

what ?

OpenStudy (fibonaccichick666):

yea, that is the problem I am having. Then on top of that, part b is to say it is only increasing in any interval containing 0 which isn't true

OpenStudy (solomonzelman):

my bad, icy...

OpenStudy (anonymous):

you can't simply take derivative and evaluate the limit because we don't even know if f(x) is differentiable at x = 0. All we know is that i is continuous at x = 0. You're supposed to show f'(0) > 0. Start from @ganeshie8's equation

OpenStudy (fibonaccichick666):

the issue sour is that the limit equals 0

OpenStudy (fibonaccichick666):

I graphed it to take a look https://www.desmos.com/calculator

OpenStudy (fibonaccichick666):

And within a neighborhood of 0, the deriv=0 like between .5,.5 or it is an itty bitty fluxuation between pos and neg

OpenStudy (fibonaccichick666):

ok I have a type o

OpenStudy (fibonaccichick666):

it's \(\color\red{x}+2x^2sin\frac{1}{x}\)

OpenStudy (fibonaccichick666):

which changes ganeshi'es and makes it work. Thanks guys, sorry for my idiocy

ganeshie8 (ganeshie8):

ugh

OpenStudy (anonymous):

:O

OpenStudy (fibonaccichick666):

I had copied the question wrong on my paper. but at least the confusion was warranted

ganeshie8 (ganeshie8):

np :) i was thinking the same, making the firs term x gives leaves a constant term so that we get some number for derivative

OpenStudy (fibonaccichick666):

yea, that's what I got, but now my issue is that part b is show that f is not increasing in any interval containing 0. But the whole function is increasing...?

OpenStudy (fibonaccichick666):

and I just quadruple checked no type-o here

OpenStudy (fibonaccichick666):

if f'>0 that implies increasing no?

OpenStudy (fibonaccichick666):

any thoughts?

OpenStudy (fibonaccichick666):

would you prefer I make a new question?

ganeshie8 (ganeshie8):

yes, we can try showing that the derivative oscillates between +ve andn -ve values as it approaches 0

OpenStudy (fibonaccichick666):

I found a similar problem, I'm just going to use that as a guide http://math.berkeley.edu/~borisp/MA104/MA104solutions3.pdf

OpenStudy (fibonaccichick666):

Thanks again Ganeshie, Sour, and Solomon

ganeshie8 (ganeshie8):

\[f'(x) = 4x\sin(1/x) - 2\cos(1/x) +1\]

ganeshie8 (ganeshie8):

consider the limit x->0 and show that the derivative function oscialletes "infinitely" between +ve and -ve values

ganeshie8 (ganeshie8):

that proves there exists no interval containing 0 in which the function is increasing (also it proves there is no interval containing 0 in which the function is decreasong)

OpenStudy (fibonaccichick666):

ok, it was a technicality I didn't understand. If it oscilates, I assumed that meant it both inc and dec

ganeshie8 (ganeshie8):

Ohk.. that pdf looks great!

ganeshie8 (ganeshie8):

stick to the definition of "increasing function"

ganeshie8 (ganeshie8):

you need to have an "interval" for which the derivative must "maintain" its sign for the entire interval

OpenStudy (fibonaccichick666):

Yea, it does it the same way you did. You go smarty pants! :)

OpenStudy (fibonaccichick666):

ohhh maintain... that makes sense now

ganeshie8 (ganeshie8):

by proving the derivative is "infinitely oscillating" around 0, we are proving that there is no "interval" around 0 in which the derivative "maintains" its sign

OpenStudy (fibonaccichick666):

That makes sense now. I did not realize that definition of increasing.

ganeshie8 (ganeshie8):